Fill in the following blanks:
15 / 18 = ☐ / 6 = 10 / ☐ = ☐ / 30 [Are these equivalent ratios?]

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

15 / 18 = (5 × 3) / (6 × 3)
= 5 / 6
5 / 6 = (5×2) / (6 × 2)
= 10 / 12
5 / 6 = (5×5) / (6 × 5)
25 / 30
Hence, 5, 12 and 25 are the numbers which come in the blanks respectively.
Yes, all are equivalent ratios.

Thanks